Using one of the laws of exponents, prove that any number raised to the 0'th power equals 1.
Answer by stanbon(75887)

Any non-zero number divided by itself is one: e.g. n/n=1 for n not equal to zero.
n^0 = n^(1-1) = n^1(n^-1)=n^1/n^1=n/n=1
Cheers,
Stan H.